OpenStudy (anonymous):

Unit 6 Rational Expression Lesson 6 Adding and subtracting Rational Expression and like denominator. Simplify n - 2 2 - n n - 2 A. 2 -n n + 2 B. -1 C. n + 2 n - 2 D. 1 my answer: C

OpenStudy (anonymous):

@Awolflover1 @jagr2713

OpenStudy (anonymous):

\[\frac{ n }{ 2-n } - \frac{ 2 }{ n-2 }\]

OpenStudy (anonymous):

Is that your question

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

it would really be helpful if you used the equation tool below fortunately @iambatman is good a mind reading

OpenStudy (anonymous):

just another superpower i guess

OpenStudy (anonymous):

I can't is not working on my side

OpenStudy (anonymous):

\[\huge \frac{ a }{ b }-\frac{ c }{ d } = \frac{ ad-bc }{ bd }\] you could use this property

OpenStudy (anonymous):

\[\frac{ n(n-2)-2(2-n) }{ (2-n)(n-2) }\]

OpenStudy (anonymous):

just to get it out there I like that username Batman

OpenStudy (anonymous):

Just start simplifying now, and thanks haha :)

OpenStudy (anonymous):

okay ur welcome

OpenStudy (anonymous):

so it will be A

OpenStudy (anonymous):

Ehh how did you manage to get that?

OpenStudy (anonymous):

because I have to cancel out the ( 2 -n) and (n-2) and that what I got

OpenStudy (anonymous):

\[\frac{ n(n-2)-2(2-n) }{ (2-n)(n-2) } \implies \frac{ n^2-2n-4+2n }{ (2-n)(n-2) }\] right?

OpenStudy (anonymous):

\[\frac{ n^2-4 }{ (2-n)(n-2) } \implies \frac{ (n-2)(n+2) }{ (2-n)(n-2) } \implies \frac{ n+2}{ 2-n }\]

OpenStudy (anonymous):

okay cool I have 3 more question it hard multitasking

OpenStudy (anonymous):

Focus on one question at a time, but if they are similar go over my work and try them out :)

OpenStudy (anonymous):

okay thanks so much